# Active Measurement: Efficient Estimation at Scale

### v1

@@ 1 Introduction
As a concrete example, consider counting birds in the high-resolution photograph in Fig. LABEL:fig:overview of a large flock of Tree Swallows in Old Lyme, CT in September, 2018. Scientists would like to estimate the population size and track it over time from photographs like this. Computer vision is an excellent tool for this task, but current workflows are poorly suited to scientific measurement. A scientist might be able to estimate the count with relatively low effort using a pre-trained object detector adapted with a few labeled examples. However, substantial additional effort is needed to obtain a model that is accurate enough for scientific measurement and gain confidence in its performance. A typical scenario would be to first tune and validate the detector’s performance using labeled image tiles. To obtain bounds on count performance, more validation is needed. After substantial effort, the scientist may find that the count is $20,358$$\pm$ $8000$ birds, which is not accurate enough to track populations over time. What can be done? Under current practice, the scientist returns to model development. For challenging tasks, it may be impossible to achieve precise enough measurements.
@@ 3 Related work
Active measurement is closely related to three existing lines of work. The form of the estimator in Eq. (1) is identical to that used by active testing [10, 16], with the significant difference that the estimand is different: active testing seeks to estimate the test loss of a fixed AI model by interactively acquiring test labels and refining a surrogate model to predict the test loss, while active measurement seeks to directly estimate a scientific measurement while iteratively refining the AI model itself. Compared to active testing, we also contribute novel weighting schemes (§ 4) and show that these reduce error, and novel approaches for variance estimation and confidence interval construction (§ 5), which was not considered in active testing.
Active testing and active measurement are both versions of adaptive importance sampling (AIS) [22, 23, 5], but modified to sample without replacement from a finite sample space, which leads to the estimator in Eq. (1), compared to the usual importance-sampling estimator of $f(s)/q(s)$ with $q$ supported on all of $\Omega$. Sampling without replacement leads to novel considerations in the selection of weights (§ 4) and in variance estimation (§ 5). The DISCount (detector-based importance sampling) method of Perez et al. [27], which builds on [20], shares the goal of counting in scientific collections using AI model predictions for the importance sampling proposal distribution. Active measurement advances on DISCount by interactively refining the AI model with acquired labels and by sampling without replacement, which both can significantly reduce estimation error.
The concept of combining an AI model with limited human effort is related to semi-supervised learning [6], where the model is trained with using a combination of labeled and unlabeled data. Most recently, prediction-powered inference (PPI) [3] was proposed to give valid statistical inference using a small amount of labeled data. Unlike PPI, which assumes data is iid and seeks to estimate a population parameter, active measurement samples data interactively and non-uniformly as aided by AI and seeks to estimate a scientific measurement on a finite dataset. In a similar spirit, active machine learning has been used to construct large datasets with crowd-sourcing [25, 17, 35]. We have a different objective of achieving accurate estimation with a small amount of labeled data.
@@ 5 Variance estimation and confidence intervals
Confidence intervals. Several practical steps remain to construct confidence intervals. The full variance estimation procedure is shown include Alg. 2. Previously, we assumed knowledge of $F(\Omega\backslash\operatorname{\mathcal{D}}_{\tau})$ for estimating the conditional variance, but in practice we use the plug-in estimator $\hat{G}_{\tau,t}=\hat{F}_{1:t-1}-F(\operatorname{\mathcal{D}}_{\tau})\approx F(\Omega\backslash\operatorname{\mathcal{D}}_{\tau})$ from step $t$ of active measurement. The full conditional variance estimator is then $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}=\sum_{\tau=1}^{t}\operatorname{\bar{\alpha}}_{\tau}^{2}\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}\approx\sum_{\tau=1}^{t}\operatorname{\bar{\alpha}}_{\tau}^{2}\operatorname{\mathrm{Var}}[F_{\tau}|\operatorname{\mathcal{D}}_{\tau}]=V^{2}_{1:t}$, which according to Prop. 5 is the appropriate quantity for the martingale central limit theorem. The naive implementation to compute $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}$ using Alg. 2 takes $\mathcal{O}(t^{2})$ time but we can reduce the complexity to $\mathcal{O}(t)$ with a streaming algorithm; see §B.2 for details. Finally, a $1-\alpha$ confidence interval is constructed as $\hat{F}_{1:t}\pm z_{\alpha/2}\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}^{1/2}$, where $z_{\alpha/2}$ is the $21-\alpha/$ 2 quantile of the standard normal distribution.
@@ 6 Experiments
We experiment with two tasks: 1) counting birds in high-resolution images of Tree Swallow flocks, and 2) counting roosting birds in weather radar images across multiple radar stations and years. The first task is analogous to common counting problems in microscopy or medical imaging, where a object detector trained on a few examples may perform reasonably well. In contrast, the second task is considerably more challenging, as it requires a custom detector—typically the result of substantial community effort involving dataset annotation, model training, and validation. For all IS-based methods, we use proposal distributions proportional to predicted counts using an object detector.
Counting birds in high-resolution images. We aim to estimate the total number of birds in two separate images—“sky" and “reeds"—each with different difficulty levels. The reeds image is more challenging due to its higher bird density and more complex background (see Fig. LABEL:fig:overview and Fig. A1). We divide the sky and reeds images into tiles of size 200$\times$200 and 160$\times$160 pixels, respectively, and manually annotate the birds in each tile using the VGG annotator [9]. This results in 925 tiles for sky and 1,426 tiles for reeds, which serve as the annotation units in our experiments. In total, the ground truth bird count is 5,682 for the sky image and 12,486 for the reeds image.
@@ 7 Results
Figure 3: Relative errors compared to $\alpha^{\mathrm{COMB}}$ weighting. Other fixed weighting strategies ($\alpha^{\mathrm{SQRT}}$, $\alpha^{\mathrm{LURE}}$) are worse, but inverse variance weighting (denoted by INV ($\gamma=0.5$)) may achieve lower error.
Figure 4: Coverage and radius (relative to the ground-truth) of CIs on the roost data for station KDLH as a function of $t$ (from 5,000 replications), built with either variance estimator from § 5. We achieve the desired coverage with narrower CIs.
@@
Figure 5: Fractional error with other baselines.
We use an experiment to illustrate the differences between active measurement and other baselines. Assume the loss function is $|f(s)-g(s)|$ for any $s\in\Omega$. Active testing will use an acquisition distribution proportional to the approximate loss to select units to label. Even if we know the ground-truth loss values, this acquisition distribution is worse than active measurement because of the mismatch between the proposal distribution and the estimand. Similarly, another estimator, motivated by PPI, is $\hat{H}_{t}=\sum_{s\in\Omega}g(s)-\frac{N}{t}\sum_{\tau=1}^{t}(g(s_{\tau})-f(s_{\tau}))$ with uniformly sampled labeled data. This estimator is also worse than active measurement because it requires a fixed detector and does not select the best labels for estimation. This is shown in Fig. 5.
@@ 8 Conclusions, Limitations, Future Work
There are several promising directions for future work. Beyond simple fine-tuning, unsupervised, semi-supervised, and transductive approaches could improve predictions across the full dataset. The current importance sampling strategy prioritizes units with high counts, but early-stage measurements could adopt more balanced sampling to better support detector training—for example, through active learning. Although model adaptation improves performance, there is a substantial computational overhead; in our experiments, we fine-tuned deep networks for detection tasks on GPUs. To enable interactivity, future systems must support fast and lightweight model updates. Approaches such as few-shot counting models and in-context learning architectures may offer practical solutions for efficient adaptation.
@@
We thank Maria C. T. D. Belotti for helpful discussions and contributing the algorithm to estimate bird counts from radar data. This work is supported in part by NSF grants #2406687, #2329927, and #2210979.
@@ A.4 Proof of proposition 4
Since $n\geq 1$, $t\geq 1$, and $n\leq t$, we can note that $u(n)>0$, $w(t)>0$, $u(n)<w(t)$, and $0.50.5\leq c_{1}(t,\lambda)\leq t-$ 0.5. This alleviates concerns over dividing by zero going forward.
@@
This is a cubic equation over $u(n)$. Looking at the coefficients, 1 and $2w(t)$ are both positive. Making a substitution and using a bound on $c_{1}$, $02w(t)^{2}-c_{1}(t,\lambda)\geq 2t+1-t+0.5=t+1.5>$ 0. Clearly the last term is negative. By Descartes’ rule of signs, the coefficients of this polynomial have 1 sign change and thus at most 1 positive real root over $u(n)$. Since $u(n)$ is monotonic over $n$ and positive for all $n\geq 1$, we see that there exists at most one $n$ such that $\frac{\partial g(n,t,\lambda)}{\partial n}=0$, for any given $t$ and $\lambda$. From this we can deduce that the equation $g(n,t,\lambda)=h(t,\lambda)$ for any function $h(t,\lambda)$ that does not depend on $n$ has at most 2 solutions over $n$.
@@ A.5 Proof of Proposition 5 (martingale CLT)
[…]
@@
[…]
@@
[…]
@@ A.7 Proof of proposition 7
The above two inequalities still hold when $t=N$, where LURE will assign infinite weights on the exact variance $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau,N}=\operatorname{\mathrm{Var}}[\hat{F}_{\tau}|D_{\tau}]$, so Var[$\operatorname{\mathrm{Var}}[\smash{\widehat{\operatorname{\mathrm{Var}}}}^{\mathrm{LURE}}_{\tau}|\operatorname{\mathcal{D}}_{\tau}]=0$. In summary, we see that $\operatorname{\mathrm{Var}}\left[\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}^{\mathrm{LURE}}|\operatorname{\mathcal{D}}_{\tau}\right]\leq\frac{E}{t-\tau+1}\cdot\min\left(1,\frac{(N-t)^{2}}{(t-\tau+1)}\right)$, for $E=\frac{\pi^{2}D(N-\tau+1)^{2}}{6}$. Thus $\operatorname{\mathrm{Var}}\left[\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}^{\mathrm{LURE}}|\operatorname{\mathcal{D}}_{\tau}\right]\lesssim\frac{1}{t-\tau+1}\cdot\min\left(1,\frac{(N-t)^{2}}{(t-\tau+1)}\right)$.

### v2

@@ 1 Introduction
As a concrete example, consider counting birds in the high-resolution photograph in Fig. LABEL:fig:overview of a large flock of Tree Swallows in Old Lyme, CT in September, 2018. Scientists would like to estimate the population size and track it over time from photographs like this [e.g., 4, 18, 7]. Computer vision is an excellent tool for this task, but current workflows are poorly suited to scientific measurement. A scientist might be able to estimate the count with relatively low effort using a pre-trained object detector adapted with a few labeled examples. However, substantial additional effort is needed to obtain a model that is accurate enough for scientific measurement and gain confidence in its performance. A typical scenario would be to first tune and validate the detector’s performance using labeled image tiles. To obtain bounds on count performance, more validation is needed. After substantial effort, the scientist may find that the count is $20,358$ $\pm$ 80008000 birds, which is not accurate enough to track populations over time. What can be done? Under current practice, the scientist returns to model development. For challenging tasks, it may be impossible to achieve precise enough measurements.
@@
Code for this paper is available at: https://github.com/cvl-umass/active-measurement.
@@ 3 Related work
Active measurement is closely related to three existing lines of work. Active testing [12, 20] interactively estimates the test loss of an AI model by sampling unlabeled points according to predicted losses from a surrogate model and then using the newly acquired labels to form an importance-sampling based estimator of the loss and update the surrogate model. The sampling strategy and form of the estimator in Eq. (1) is identical to that used by active testing, with the significant difference that the estimand is different: active measurement seeks to directly estimate a scientific measurement while iteratively refining the AI model itself, while active testing estimates a test loss and refines a surrogate model. Compared to active testing, we also contribute novel weighting schemes (§ 4) and show that these reduce error, and novel approaches for variance estimation and confidence interval construction (§ 5), which was not considered in active testing.
Active testing and active measurement are both versions of adaptive importance sampling (AIS) [27, 28, 6]. AIS improves an importance-sampling estimator $f(s)/q(s)$ by iteratively refining the proposal $q$ and forming a weighted combination of the estimators made with different proposal distributions. Active measurement modifies AIS to sample without replacement from a finite sample space, which leads to the estimator in Eq. (1), compared to the usual importance-sampling estimator supported on all of $\Omega$. Sampling without replacement leads to novel considerations in the selection of weights (§ 4) and in variance estimation (§ 5). The DISCount (detector-based importance sampling) method of Perez et al. [32], which builds on [25], estimates the total counts in scientific collections using AI model predictions for the importance sampling proposal distribution. Active measurement shares the same goal, and advances on DISCount by interactively refining the AI model with acquired labels and by sampling without replacement, which both can significantly reduce estimation error.
The concept of combining an AI model with limited human effort is related to semi-supervised learning [8], where the model is trained with using a combination of labeled and unlabeled data. Most recently, prediction-powered inference (PPI) [3] was proposed to give valid statistical inference using a small amount of labeled data. Unlike PPI, which assumes data is iid and seeks to estimate a population parameter, active measurement samples data interactively and non-uniformly as aided by AI and seeks to estimate a scientific measurement on a finite dataset. We show in the experiments that active measurement outperforms a PPI-motivated baseline. In a similar spirit, active machine learning has been used to construct large datasets with crowd-sourcing [30, 21, 40]. We have a different objective of achieving accurate estimation with a small amount of labeled data.
@@ 5 Variance estimation and confidence intervals
Confidence intervals. Several practical steps remain to construct confidence intervals. The full variance estimation procedure is shown include Alg. 2. Previously, we assumed knowledge of $F(\Omega\backslash\operatorname{\mathcal{D}}_{\tau})$ for estimating the conditional variance, but in practice we use the plug-in estimator $\hat{G}_{\tau,t}=\hat{F}_{1:t-1}-F(\operatorname{\mathcal{D}}_{\tau})\approx F(\Omega\backslash\operatorname{\mathcal{D}}_{\tau})$ from step $t$ of active measurement. The full conditional variance estimator is then $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}=\sum_{\tau=1}^{t}\operatorname{\bar{\alpha}}_{\tau}^{2}\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}\approx\sum_{\tau=1}^{t}\operatorname{\bar{\alpha}}_{\tau}^{2}\operatorname{\mathrm{Var}}[F_{\tau}|\operatorname{\mathcal{D}}_{\tau}]=V^{2}_{1:t}$, which according to Prop. 5 is the appropriate quantity for the martingale central limit theorem. The naive implementation to compute $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}$ using Alg. 2 takes $\mathcal{O}(t^{2})$ time but we can reduce the complexity to $\mathcal{O}(t)$ with a streaming algorithm; see §B.2 for details. Finally, a $1-\alpha$ confidence interval is constructed as $\hat{F}_{1:t}\pm z_{\alpha/2}\smash{\widehat{\operatorname{\mathrm{Var}}}}_{1:t}^{1/2}$, where $z_{\alpha/2}$ is the $1-\alpha/2$ quantile of the standard normal distribution.
@@ 6 Experiments
We mainly experiment with two tasks: 1) counting birds in high-resolution images of Tree Swallow flocks, and 2) counting roosting birds in weather radar images across multiple radar stations and years. The first task is analogous to common counting problems in microscopy or medical imaging, where an object detector trained on a few examples may perform reasonably well. In contrast, the second task is considerably more challenging, as it requires a custom detector—typically the result of substantial community effort involving dataset annotation, model training, and validation. To demonstrate domain generality, we also performed experiments on malaria-infected cell counting and damaged-building counting from satellite images. For all IS-based methods, we use proposal distributions proportional to predicted counts using an object detector.
Counting birds in high-resolution images. We aim to estimate the total number of birds in two separate images—“sky" and “reeds"—each with different difficulty levels. The reeds image is more challenging due to its higher bird density and more complex background (see Fig. LABEL:fig:overview and Fig. A1). We divide the sky and reeds images into tiles of size $\times200$ and $\times160$ pixels, respectively, and manually annotate the birds in each tile using the VGG annotator [11]. This results in 925 tiles for sky and 1,426 tiles for reeds, which serve as the annotation units in our experiments. In total, the ground truth bird count is 5,682 for the sky image and 12,486 for the reeds image.
@@
Counting malaria-infected cells.
The Malaria Cell dataset (image set BBBC041v1, available from the Broad Bioimage Benchmark Collection [24]) comprises 1,364 images (about 80,000 cells). This dataset contains microscopy images of blood smears used for detecting and classifying malaria parasites, aimed at enabling the development and evaluation of automated methods for parasite detection, quantification, and stage classification. For this evaluation, we focus on counting the number of infected cells, which are around 5% of the dataset. We use the same settings as our sky and reeds image experiments. For our initial model we finetune the default Faster R-CNN network on three randomly selected cell images.
Counting damaged buildings.
For damaged building detection we focus on the Palu Tsunami subset of xBD [15], which contains 113 satellite images of the shoreline before and after the Palu Tsunami. We count the number of damaged buildings, which corresponds to a label of “major-damage” or “destroyed”. Only the post disaster images are given to the model. We use the same hyperparameters as before. The initial model is trained on 5 randomly selected images.
@@ 7 Results
Figure 3: Relative errors compared to $\alpha^{\mathrm{COMB}}$ weighting. Other fixed weighting strategies ($\alpha^{\mathrm{SQRT}}$, $\alpha^{\mathrm{LURE}}$) are worse, but inverse variance weighting (denoted by INV ($\gamma=0.5$)) may achieve lower error. which match our theories
Figure 4: Coverage and radius (relative to the ground-truth) of CIs on the roost data for station KDLH as a function of $t$ (from 5,000 replications), built with either variance estimators from § 5. The left panel uses the $\smash{\widehat{\text{Var}}}^{\text{cond}}_{1:t}$ estimator. We achieve the desired coverage with narrower CIs.
@@
Figure 5: Fractional error compared with other baselines ($\hat{H}$ is motivated by PPI).
Figure 6: Estimation error on additional datasets, averaged over 1,000 end-to-end trials.
We use an experiment to illustrate the differences between active measurement and other baselines. Assume the loss function is |f(s)g(s)||f(s)-g(s)| for any $s\in\Omega$. Active testing will use an acquisition distribution proportional to the approximate loss to select units to label. Even if we know the ground-truth loss values, this acquisition distribution is worse than active measurement because of the mismatch between the proposal distribution and the estimand. Similarly, another estimator, motivated by PPI, is $\hat{H}_{t}=\sum_{s\in\Omega}g(s)-\frac{N}{t}\sum_{\tau=1}^{t}(g(s_{\tau})-f(s_{\tau}))$ with uniformly sampled labeled data. This estimator is also worse than active measurement because it requires a fixed detector and does not select the best labels for estimation. This is shown in Fig. 6.
Additional results.
Additional experimental results on counting malaria cells and counting damaged buildings are in Fig. 6. Here we run active measurement with end-to-end settings and average over 1,000 trials. From the first figure, we observe trends similar to those seen in the image and radar experiments for malaria cell counting. DIS consistently outperforms MC, while AIS shows an early reduction in error as the model improves. WOR also demonstrates steady improvement, with its impact becoming more pronounced after roughly 10% of the data is labeled. The results on the second dataset follow similar trends: WOR provides the most benefit when a larger fraction is labeled, while AIS is more helpful early on. We also note that our performance is lower than the numbers reported in DISCount [32], as they use a detector specifically trained for damaged building detection, whereas we use a simple ImageNet-pretrained backbone fine-tuned on just 5 satellite images. These results show that active measurement is capable of improving scientific estimation tasks across different domains.
@@ 8 Conclusions, Limitations, Future Work
There are several promising directions for future work. Beyond simple fine-tuning, unsupervised, semi-supervised, and transductive approaches could improve predictions across the full dataset. The current importance sampling strategy prioritizes units with high counts, but early-stage measurements could adopt more balanced sampling to better support detector training—for example, through active learning. Although model adaptation improves performance, there is a substantial computational overhead; in our experiments, we fine-tuned deep networks for detection tasks on GPUs. To enable interactivity, future systems must support fast and lightweight model updates. Approaches such as few-shot counting models and in-context learning architectures may offer practical solutions for efficient adaptation. Another promising opportunity is to consider correlation among images. As an example, in the reeds image we have observed spatial auto-correlation in the residuals of the AI model. Fitting a Gaussian process can exploit this spatial structure to improve predictions for unlabeled tiles.
@@
We thank Maria C. T. D. Belotti for helpful discussions and contributing the algorithm to estimate bird counts from radar data. We also thank Jody Dole for kindly granting permission to use the sky and reeds photographs featured in our main experiment. This work is supported in part by NSF grants #2504073, #2406687, #2329927, and #2210979.
@@ A.4 Proof of proposition 4
Since $n\geq 1$, $t\geq 1$, and $n\leq t$, we can note that $u(n)>0$, $w(t)>0$, $u(n)<w(t)$, and $0.5\leq c_{1}(t,\lambda)\leq t-0.5$. This alleviates concerns over dividing by zero going forward.
@@
This is a cubic equation over $u(n)$. Looking at the coefficients, 1 and $2w(t)$ are both positive. Making a substitution and using a bound on $c_{1}$, $2w(t)^{2}-c_{1}(t,\lambda)\geq 2t+1-t+0.5=t+1.5>0$. Clearly the last term is negative. By Descartes’ rule of signs, the coefficients of this polynomial have 1 sign change and thus at most 1 positive real root over $u(n)$. Since $u(n)$ is monotonic over $n$ and positive for all $n\geq 1$, we see that there exists at most one $n$ such that $\frac{\partial g(n,t,\lambda)}{\partial n}=0$, for any given $t$ and $\lambda$. From this we can deduce that the equation $g(n,t,\lambda)=h(t,\lambda)$ for any function $h(t,\lambda)$ that does not depend on $n$ has at most 2 solutions over $n$.
@@ A.5 Proof of Proposition 5 (martingale CLT)
Assume that (1) $\sum_{\tau=1}^{T_{N}}\operatorname{\tilde{\alpha}}_{\tau}^{2}\operatorname{\mathrm{Var}}[\hat{F}_{\tau}|\operatorname{\mathcal{D}}_{\tau}]\to\eta^{2}$ as $N\to\infty$; (2) $P(\eta^{2}<\infty)=1$ and $P(\eta^{2}>0)=1$; (3) $\sum_{\tau=1}^{T_{N}}\operatorname{\mathbb{E}}[X_{N,\tau}^{2}\mathbb{I}(|X_{N,\tau}|>\epsilon)|\operatorname{\mathcal{D}}_{\tau}]\xrightarrow{p}0$ for any $\epsilon>0$. We have two central limit theorems
$\frac{S_{N,T_{N}}}{\sqrt{\sum_{\tau=1}^{T_{N}}X_{N,\tau}^{2}}}\xrightarrow{D}\mathcal{N}(0,1),$ (11)
$\frac{S_{N,T_{N}}}{\sqrt{\sum_{\tau=1}^{T_{N}}\operatorname{\tilde{\alpha}}_{\tau}^{2}\operatorname{\mathrm{Var}}[\hat{F}_{\tau}|\operatorname{\mathcal{D}}_{\tau}]}}\xrightarrow{D}\mathcal{N}(0,1).$ (12)
@@
$\frac{S_{N,T_{N}}}{U_{N,T_{N}}}\xrightarrow{D}\mathcal{N}(0,1),$
$\frac{S_{N,T_{N}}}{V_{N,T_{N}}}\xrightarrow{D}\mathcal{N}(0,1),$
@@
$\frac{\hat{F}_{1:T_{N}}-F(\Omega)}{\sqrt{\sum_{\tau=1}^{T_{N}}\operatorname{\bar{\alpha}}_{\tau}^{2}(\hat{F}_{\tau}-F(\Omega))^{2}}}\xrightarrow{D}\mathcal{N}(0,1),$ (13)
$\frac{\hat{F}_{1:T_{N}}-F(\Omega)}{\sqrt{\sum_{\tau=1}^{T_{N}}\operatorname{\bar{\alpha}}_{\tau}^{2}\operatorname{\mathrm{Var}}[\hat{F}_{\tau}|\operatorname{\mathcal{D}}_{\tau}]}}\xrightarrow{D}\mathcal{N}(0,1).$ (14)
@@ A.7 Proof of proposition 7
The above two inequalities still hold when $t=N$, where LURE will assign infinite weights on the exact variance $\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau,N}=\operatorname{\mathrm{Var}}[\hat{F}_{\tau}|D_{\tau}]$, so $\operatorname{\mathrm{Var}}[\smash{\widehat{\operatorname{\mathrm{Var}}}}^{\mathrm{LURE}}_{\tau}|\operatorname{\mathcal{D}}_{\tau}]=0$. In summary, we see that $\operatorname{\mathrm{Var}}\left[\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}^{\mathrm{LURE}}|\operatorname{\mathcal{D}}_{\tau}\right]\leq\frac{E}{t-\tau+1}\cdot\min\left(1,\frac{(N-t)^{2}}{(t-\tau+1)}\right)$, for $E=\frac{\pi^{2}D(N-\tau+1)^{2}}{6}$. Thus $\operatorname{\mathrm{Var}}\left[\smash{\widehat{\operatorname{\mathrm{Var}}}}_{\tau}^{\mathrm{LURE}}|\operatorname{\mathcal{D}}_{\tau}\right]\lesssim\frac{1}{t-\tau+1}\cdot\min\left(1,\frac{(N-t)^{2}}{(t-\tau+1)}\right)$.